a)     Determine the Fixed Cost, Average Variable cost, and Marginal cost functions for a firm whose total cost function is defined by

 

C = 36x-10x^2+2x^3 


1
Expert's answer
2021-11-11T13:26:02-0500

"Marginalcost(x)=c'(x)=VC'(x)=(36x-10x^2+2x^3)'=36-20x+6x^2"

TC=FC+VC

"TC=36x-10x^2+2x^3"

FC=lim+C(x)

x→0

FC=0

"AVC=\frac{VC}{Q}"

Q=x

"AVC=\frac{VC}{Q}=\frac{36x-10x^2+2x^3}{x}=36-10x+2x^2"